OpenStudy (anonymous):

If f(x) = ∣(x^2 − 6)(x^2 + 2)∣, how many numbers in the interval 1 ≤ x ≤ 2 satisfy the conclusion of the mean value theorem? a. none b. one c. two d. three

OpenStudy (anonymous):

@Abhisar @Hero @iambatman @Michele_Laino @satellite73 @TheSmartOne @Zale101

OpenStudy (michele_laino):

please note that within the interval (1,2) x^2-6 is negative

OpenStudy (anonymous):

@Michele_Laino what do you mean?

OpenStudy (michele_laino):

I meant that \[f(x) = \left( {6 - {x^2}} \right)\left( {{x^2} + 2} \right)\] when x is in [1,2]

OpenStudy (michele_laino):

we have to solve this equation: \[4{x^3} - 8x - 3 = 0\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

i got 3 is that right @Michele_Laino